OpenStudy (anonymous):

Derivative: If f(x)=√24x, find f′(x).

OpenStudy (anonymous):

what is under the radical, 24, or 24x ?

OpenStudy (anonymous):

24x is under the radical sign.

OpenStudy (anonymous):

then think of it as \[\sqrt{24}\sqrt{x}\] leave the constant there, and the derivative of \(\sqrt{x}\) is \(\frac{1}{2\sqrt{x}}\)

OpenStudy (anonymous):

Can I use the product rule since they are multiplying?

OpenStudy (anonymous):

you do not use the product rule for constants \[\frac{d}{dx}[cf(x)]=cf'(x)\]

OpenStudy (anonymous):

Thank you for your help.